OpenStudy (anonymous):

guitar list price is 1299.90 with a chain discount of 5/3/2. What would be the trade discount amount? what?

OpenStudy (anonymous):

5/3/2 = 10/3 , so its looks strange to me. Are u sure that everything is correct

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

10/3 = 3 1/3, so it is 3 and one third of something. I guess it cant be the prize, becouse they would pay to you for buying, :)